OpenStudy (anonymous):

Daniel cold buy 8 apples and 5 peaches with 8$ he could buy 16 such apples with the same amount of money. if he decided to buy peaches only,how many peaches cold he buy with $100 ?

OpenStudy (anonymous):

Could*

OpenStudy (anonymous):

8x+5y=8 16x=8 8x=4 4+5y=8 5y=4 y=4/5 100/(4/5)=5*100/4 =125